OpenStudy (anonymous):

Find an exact value. tan 75°

OpenStudy (anonymous):

3.732050807568877.......

OpenStudy (unklerhaukus):

notice that 75=30+45

OpenStudy (unklerhaukus):

\[\tan(u+v)=\frac{\tan(u)+\tan(v)}{1-\tan(u)\tan(v)}\]

OpenStudy (unklerhaukus):

you'll have to know [or be able to determine] the exact value of tan(30) and tan(45)